OpenStudy (tay0913):

WAT is the answer to 6x-3=2(3x+4)

OpenStudy (tay0913):

I need help

OpenStudy (tay0913):

Plz

OpenStudy (faiqraees):

equation is invalid

OpenStudy (tay0913):

That's what it says on my paper

OpenStudy (tay0913):

But thx u

OpenStudy (faiqraees):

they must be testing you to see if you can see the invalidity in the equation.

OpenStudy (faiqraees):

\[6x-3 = 2(3x+4) \] \[6x-3= 6x+8\] Cancel 6x on both sides and you're left with -3=4 which isn't possible.

OpenStudy (tay0913):

So wats the answer